The time it takes to travel to a location varies inversely to the speed traveled. It takes 4 hours driving at an average speed of 55 mile per hour to reach a location. To the nearest tenth of an hour, how long will it take to reach the same location driving at an average speed of 50 miles per hour?
Answer by checkley71(8403)

DISTANCE=RATE*TIME
D=55*4
D=50T
SEEING AS THE DISTANCES ARE THE SAME WE HAVE:
55*4=50T
220=50T
T=220/50
T=4.4 HOURS FOR THE SLOWER DRIVER.